# biblio.bib

### v1

@@ 2.2 Role of the anisotropy (ZZ-term interaction)
Once the energy bands of the exchange Hamiltonian ^$\hat{\mathcal{H}}_{0}$ are computed, we can consider the effect of the anisotropy/interaction term in the Hamiltonian (1):
@@
Since the levels are degenerate, we should in principle compute (at first order) the eigenvalues of the $\hat{\mathcal{W}}$ perturbation restricted to the considered eigenspace. However, it can be easily checked by inspection that $\hat{\mathcal{W}}$ is already expressed in a diagonal form, due to the definitions of the Pauli basis $\mathcal{P}$. Therefore, we only need a restricted number of matrix elements. For any form of the $\Delta_{ij}$ coefficients one can cast the action of the perturbation operator on a given state in the computation basis $|\boldsymbol{\sigma}_{q}\rangle$, with $\boldsymbol{\sigma}_{q}$ expressing the spin polarizations vector in each site, the index $q$ denoting the specific Pauli string. We thus obtain:
@@
where sgn(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denotes the sign function. We can note how the shifts in energy depend on the pairwise products of the spins’ directions in the basis vector. In order to simplify our analytical computation, we consider now nearest-neighbor interaction, in other words $\Delta_{ij}=\Delta_{0}$ if dist$(i,j)=1$ and vanishing otherwise. Under this condition, the energy shifts read:
[…]
@@
Without loss of generality, we can hence turn to examine in detail the case of nearest-neighbor interaction for the perturbation. The shift reflects the fact that if there is a total number of $L$ couples (with PBC) and $b$ is the number of pairs with opposite spin directions in a given Pauli string, the sum (10) corresponds to the number of consecutive pairs with spins pointing in the same direction after subtracting those with opposite spin, that amounts to $(L-b)-b=L-2b$. By way of example, let us compare the resulting effect for the cases of one and two excitations. In the former case, there are only two out of the $L$ couples of spins that have opposite directions. Thus, the summation in (10) amounts to $L-4$ for each vector state of the restricted Pauli basis $\mathcal{P}_{1}$. This means that the shift is the same for both levels: importantly the perturbation has no overall effect on the local $z$-magnetization frequency. In the latter case, one can distinguish two possibilities for the Pauli strings, as depicted in table 1. According to the considered vector, one can get different values for the summation in eq. (10) and hence for the energy shift.
@@
Interestingly, this result can also be linked to the description of information processing properties of the out-of-equilibrium XXZ network with all-to-all connectivity analyzed in [YagoMalo2023, yagomalo2024review]. There, an ideal-observer analysis of the power spectra had shown that the uncertainty with which the quantum network encodes the number of frequency peaks - and therefore the number of excitations - linearly scales with the number itself. In neuroscience, this is the content of Weber’s law describing the perception of the number of items in a space, or numerosity perception. This phenomenology is experimentally observed under very general conditions, no matter the perception channel (i.e. visual or auditory), nor the way the items are presented [burr2008visual]. Importantly, neuroscience experiments indicate that the perception of number is interconnected to the perception of time and space intervals [Morrone:2007].
The overall bottom line is that even in the XXZ quantum network the emerging dynamical frequencies - or else the numerosity counting itself- comes from the number of consecutive gaps between the original levels. In fact, up to perturbative levels, anisotropy is found to just inject an uncertainty due to degeneracy lifting and leading to consequent spreading of the bands, the latter depending itself on the number of excitations introduced.
With this physical understanding at hand, we can now turn our attention to the transport properties of more general quantum networks, where we release either the constraint of constant coupling intensity of the links in the regular network, or the regular architecture itself by considering complex quantum networks. This is the subject of the next Section 3, where we explore the extent to which the space and time memory-preserving property of the regular- XXZ network excitations can be engineered in more general networks.
@@ 3 Transport properties in generalized quantum networks
So far we have exploited the concept of cooperative shielding to build intuition and theoretical understanding about the out-of-equilibrium dynamical behavior of the closed quantum spin network governed by the XXZ Hamiltonian, for a network with all-to-all connectivity architecture of equally weighted links. For this case, we could resort to semi-analytical means. In this section, we start from the understanding developed in Section 2 to numerically investigate the role played by the interaction graph in a complex quantum system architecture. For simplicity, we stick to case of one single excitation, having developed in Section 2 an intuition for multiple excitations: we will come back to this point in the concluding Section 4. As an example for complex quantum network we choose the small-world architecture: besides being per se relevant for diverse problems, we are here interested to its property of well interpolating between the cases of regular and random graphs after tuning only two parameters. We analyze the underlying physics by specializing our analysis to two relevant quantities, a localization quantifier called Inverse Participation Ratio (IPR) (see Appendix 1) and the relative distance $\Delta_{\text{IPR}}$ introduced to measure how much the IPRs of the small-world graphs depart from the all-to-all networks (see below).
@@ 3.1 Small-world networks
A number of real-life networks relevant to diverse phenomena, ranging from biology to social sciences, are characterized by a high level of clustering and, at the same time, by a low average distance between the vertices: this is the so-called “small-world phenomenon” that illustrates e.g. social networks and the World Wide Web [watts1998smallworld].
@@
Such a network can be built from a regular-lattice graph where each node is linked to its first $2k$ neighbors, and then introducing a rewiring probability $\beta$ that one edge connects the starting node to a different one. The parameter $2k$ drives the degree of network complexity. The parameter $\beta$ governs the level of graph’s randomness, bridging between the regular-lattice case with $\beta=0$ and the fully random graph with $\beta=1$. The all-to-all connected graph is obtained in the limit $\beta=0$ and $12k/(L-1)=$ 1. The construction mechanism is illustrated in Fig. 4$(a)$.
@@
We complete this introductory section by recalling the main qualifiers and quantifiers for the networks architectures considered in complex-network theory, that will be used in the following. For details, we refer to the Appendix. One first connectivity quantifier is the network’s degree, that is the number of edges linking one node to others. The coupling range is quantified by the topological distance distij, counting the number of links connecting two sites in their shortest path: a global distance quantifier $\mathcal{D}$ can then be usefully introduced, as the average of the topological distance over all the pairs of nodes. Finally, the tendency to aggregate is quantified by the clustering $\mathcal{C}$, that is the average proportion of total triangles (meaning closed loops of three vertices) within a given neighborhood, as related to the maximum number of allowed connections according to the local degree. All-to-all connected graphs are the limiting case of low topological distance, i.e. tending to 1, and high clustering coefficient, i.e. tending to 1; in the opposite limiting case of nearest-neighbor networks both local and global clustering are 0, since no triangles between nodes can be observed, while average distance linearly scales as the number of sites $L$.
@@ 3.2 Inverse Participation Ratio
Figure 4: Construction of small-world graphs and their transport properties. $(a)$ Small-world network construction (see text). Top to bottom: a regular graph with $L=7$ sites and $k=2$ nearest-neighbor connections, and different rewired links (green and orange) with respect to the original graph, according to the value of the probability $\beta$. $(b)-(c)$ Examples of transport for one quenched excitation along the $z$-axis in two different networks with $L=33$ sites, shown in the corresponding insets with the initial site highlighted in red (sub-figure $(b)$ has initial $k=1$ and initial site 1 with degree 1, while $(c)$ has $k=3$ and initial site 13 with degree 6). The average local $z$-magnetization $\bar{\langle\mathcal{S}_{i}^{z}\rangle}$ time signal is displayed as a function of the dimensionless time $tJ/2$ (horizontal axis), in sites (vertical axis) having the same topological distance distij from the site where the excitation has been injected. $(d)$ Time evolution of the Inverse Participation Ratio (IPR) for one excitation injected in different initial sites of a network with small rewiring probability $\beta$ and large $k$ values, for the different connectivity levels in the legend as measured by the values of the degree, i.e. the number of edges, related to $(L-1)$: values closer to 1 tend towards an all-to-all connected graph (light-blue curve). Parameters: $L=33$, $22k=L-$ 2, $\beta=0.1$. $(e)$ IPR distance (12) between all-to-all and small-world networks graphs as a function of the initial site’s degree related to $L-1$ at increasing times as in the legend.
The transport properties of one excitation in such networks can be characterized by means of a de-localization quantifier, the Inverse Participation Ratio (IPR)
@@
defined as the sum of the squared trace elements of the density matrix $\hat{\mathcal{\varrho}}$ over single-excitation localized states $\{|s\rangle\}=\{|\downarrow\downarrow\dots\downarrow\uparrow\downarrow\dots\downarrow\rangle\}$, where the $\uparrow$ in the previous Pauli string is a spin up (along $z$) located at site $s$; this basis has dimension $L$ and constitutes one of the subspaces of the total $z$-magnetization operator ^$\hat{\mathcal{M}}_{z}$ with which $\hat{\mathcal{H}}$ can be block-diagonalized. While we refer to the Appendix for a complete definition, we here highlight that by definition the IPR values range from $1/L$ for the complete delocalized state, i.e. infinite-temperature case, to $1$ for the perfect localization on a single eigenstate. If needed, generalizations of this quantifiers to the case of multiple excitations can be introduced, as in [Civolani2024].
We begin by examining the rate of local transport in Fig. 4$(b-c)$ showing two examples for the time-evolution of the local $z$-magnetization signals according to different values of the topological distance distij for two different degrees of connectivity. We observe that in the sparsely connected case $(b)$ the initial excitation takes longer time to spread across the network with respect to more connected case $(c)$, leading to the clear emergence of a light-cone for the spreading of the excitation. As the number of connections increases $(c)$, the de-localization of the first excitation is distributed between all sites more or less evenly: therefore, the spreading occurs at very short times with and the cone structure is harder to identify. Thus, it is interesting to study increasing connectivities close to the limit of all-to-all connected networks. The results are illustrated in Fig. 4$(d)$, where we consider a network with a given $k=L-2$ parameter, closest to all-to-all case but not-fully connected, and impose a non-zero rewiring probability providing the sites with different degrees, i.e. different number of links with other elements. Then, we compute the IPR by picking different initial sites according to their degree. We observe that the IPR time-evolution behavior for the all-to-all case soon breaks into the general small-world system behavior. Here, the degeneracy lifting due to the symmetry reduction makes the IPR decreasing over time towards full delocalization (yellow, orange, and dark blue curves), instead of oscillating around the localization condition (light blue and green curves). We notice the intermediate case displayed by the purple curve, that seems to herald the passage from localization to delocalization with IPR oscillations that increase in amplitude and slow down in time.
We quantity the departure of small-world behavior from the all-to-all connectivity case in Fig. 4$(d)$, where we introduce an indicator for the difference the distance:
$\Delta_{\text{IPR}}(t)=\frac{|\text{IPR}_{\text{a.a.}}(t)-\text{IPR}_{\text{s.w.}}(t)|}{\text{IPR}_{\text{a.a.}}(t)}\,.$ (12)
Different dots colors for $\Delta_{\text{IPR}}$ refer to values computed at different time instants $tJ/2$ in units of $2\pi/\omega$, $\omega$ being determined as in Eq. 3 for the quantum numbers $l=L/2$ and $m=-L/2+1$. Each point represents the average computed over initial sites having the same degree in the same network as in Fig. 4$(d)$. As expected, we observe an increasing difference while the network’s degree departs from $L-1$ (degree/$(L-1)<1$) corresponding to all-to-all connectivity, the difference being enhanced at each degree value with increasing time.
@@
While these results are intuitively sound, as in Section 2 we search for analytical understanding. Since the analytical machinery that we established for regular networks seems to not apply for irregular graphs, we seek for alternatives. It turns out that this is possible at least for the short-time IPR behavior, by means of the so-called cascade model introduced in [borgonovi2019ipr]: in essence, this monitors the initial state $|s(0)\rangle$ coupling with other basis elements at different moments $|s(t)\rangle$.
@@
The underlying idea is as follows, see Fig. 5$(a)$ for the concept. At $t=0$, only the subset of state vectors $\mathcal{F}_{0}$ containing $|s\rangle$ participates in the temporal evolution. Immediately after, to first-order approximation the temporal evolution operator connects the initial state to the basis vectors $|f_{i}\rangle$ for which the matrix elements $\langle f_{i}|\hat{\mathcal{H}}|s\rangle$ are non-vanishing, thus extending the subset of involved states to $\mathcal{F}_{1}$. More precisely, the subset of states grows as
@@
Figure 5: Understanding transport in highly-connected graphs at early times. $(a)$ Concept of the cascade model (see text). Top: an initial excitation (red circle) in a generic network (black edges) is located at site 1, as denoted by the state $|s=1\rangle=|\uparrow\downarrow\dots\downarrow\rangle$ and the red circle opposed to the blue ones. Bottom: after a short temporal evolution (black arrow) the initial excitation has spread (gradient-colored circles) across site 1’s first neighbors (orange edges), such that the state is a superposition of $|s=1\rangle$ and states $|f_{2}\rangle=|\downarrow\uparrow\downarrow\dots\downarrow\rangle$, $|f_{4}\rangle$, $|f_{5}\rangle$ and $|f_{7}\rangle$. $(b)-(c)$ Comparison of the IPR’s time evolution as computed from the cascade model (orange curves) and from a numerical simulation (blue curves) for the small-world network depicted in the inset with $L=33$ nodes and excitation injected in different sites (red dot). The complete delocalization limit $1/L$ is shown with a dashed line. $(b)$ Case with low average distance between sites (network parameters: $k=1,\beta=0.31$). $(c)$ Case with large average distance between sites (network parameters: $k=2,\beta=0.88$). $(d)$ IPR early-time evolution for a fully-connected network with $L=33$ nodes and links intensity that changes with physical distance according to a power law with coefficient $\alpha$ as in the legend, $J_{ij}\propto\text{dist}(i,j)^{-\alpha}$, where dist$(i,j)=$min$(|i-j|,L-|i-j|)$ for the PBCs considered in all these cases. The predictions of the cascade model for the corresponding $\alpha$ are represented with same colors, but dashed lines.
@@
Borgonovi et al. have found a transparent expression for the evolution of IPR at short times, in the approximation where any back-flow to the initial site can be neglected [borgonovi2019ipr]. This is:
@@
where the sum is performed over the first-neighbor sites of the initial one, and $\mathcal{N}_{1}$ is the number of such nodes plus one (referring to vertex $s$). Let us better understand eq. 14: as the survival probability $P_{s}$ decreases because of the excitation spreading, the initial site couples with its first neighbors in the interaction topology. In other words, the IPR evolution is strictly dependent on the graph upon which it lives, through its adjacency matrix $A_{ij}$ describing the existing connections and being proportional to the XX Hamiltonian itself, when written in the restricted basis of Pauli vectors having one excitation. Notice that $A_{ij}=1$ if sites $i$ and $j$ are linked, else $A_{ij}=0$. Under a different perspective, the degree of the initial node (i.e. the number of first neighbors in the graph) governs the transport onset at short times.
@@
A summary of the analytical results is presented in Fig. 5. In Fig. 5$(b)-(c)$, we compare the simulated (blue curves) IPR time behavior with the predictions of the cascade model (orange curves) for two complex network graphs that are highly $(b)$ and sparsely $(c)$ connected, respectively. For earlier times we find a good agreement between the IPR simulations and the predictions provided by the cascade model. In fact, this is able to predict the sharp deviation from the all-to-all case, heralding the breaking of the spectral degeneracy and the halt of cooperative shielding. The later times behavior from the cascade model misses the bumps that are likely due to the backflow processes, that are neglected in the model approximations.
A different way to break space regularity in networks is to consider exchange coupling that vary with physical distance, specifically characterized by a power law, i.e. $J_{ij}\propto\text{dist}(i,j)^{-\alpha}$ in Eq. (1) which in the case of PBCs is equivalent to min$(|i-j|,L-|i-j|)$. We thus take advantage of the cascade model to check on the excitation localization properties for this topology setting. This is illustrated in Fig. 5$(d)$, where we compare the cascade model with the IPR evolution at early times for two distinct values of $|\alpha|$. We observe good agreement until $tJ/2$ times corresponding to the tunneling time to second neighbors. The regular lattice with variable distance coupling soon deviates from the all-to-all graph, the cascade model accurately predicting this deviation. Overall then, we can conclude that space regularity, at least on a local scale and already at early times, is paramount for the emergence of highly-degenerate eigenspaces that induce the trapping of excitations via cooperative shielding. Therefore, any deviation from such space regularity soon leads to delocalization of injected excitations. The question then arises about what happens on a global scale.

### v2

@@ 2.2 Role of the anisotropy (ZZ-term interaction)
Once the energy bands of the exchange Hamiltonian $\hat{\mathcal{H}}_{0}$ are computed, we can consider the effect of the anisotropy/interaction term in the Hamiltonian (1):
@@
Since the levels are degenerate, we should in principle compute (at first order) the eigenvalues of the $\hat{\mathcal{W}}$ perturbation restricted to the considered eigenspace. However, it can be easily checked by inspection that $\hat{\mathcal{W}}$ is already expressed in a diagonal form, due to the definitions of the computational basis $\mathcal{P}$. Therefore, we only need a restricted number of matrix elements. For any form of the $\Delta_{ij}$ coefficients one can cast the action of the perturbation operator on a given state in the computational basis $|\boldsymbol{\sigma}_{q}\rangle$, with $\boldsymbol{\sigma}_{q}$ expressing the spin polarizations vector in each site, the index $q$ denoting the specific Pauli string. We thus obtain:
@@
where:
$\text{sgn}(i,q)$ $=\begin{cases}+1\quad\sigma_{i}^{(q)}=\uparrow\\ -1\quad\sigma_{i}^{(q)}=\downarrow\\ \end{cases}.$
We can note how the shifts in energy depend on the pairwise products of the spins’ directions in the basis vector. In order to simplify our analytical computation, we consider now nearest-neighbor interaction, in other words $\Delta_{ij}=\Delta_{0}$ if dist(i,j)=1(i,j)=1 and vanishing otherwise. Under this condition, the energy shifts read:
$\langle\boldsymbol{\sigma}_{q}|\hat{\mathcal{W}}|\boldsymbol{\sigma}_{q}\rangle=\frac{\Delta_{0}}{4}\sum_{i=1}^{L}\text{sgn}(i,q)\text{sgn}(i+1,q)\quad,\quad L+1\equiv 1\,\text{(PBC)}\,.$ (10)
Note that PBC stands for Periodic Boundary Conditions.
@@
Without loss of generality, we can hence turn to examine in detail the case of nearest-neighbor interaction for the perturbation. The shift reflects the fact that if there is a total number of $L$ couples (with PBC) and $b$ is the number of pairs with opposite spin directions in a given Pauli string, the sum (10) corresponds to the number of consecutive pairs with spins pointing in the same direction after subtracting those with opposite spin, that amounts to (Lb)b=L2b(L-b)-b=L-2b. By way of example, let us compare the resulting effect for the cases of one and two excitations. In the former case, there are only two out of the $L$ couples of spins that have opposite directions. Thus, the summation in (10) amounts to L4L-4 for each vector state of the restricted computational basis $\mathcal{P}_{1}$. This means that the shift is the same for both levels: importantly the perturbation has no overall effect on the local $z$-magnetization frequency. In the latter case, one can distinguish two possibilities for the Pauli strings, as depicted in table 1. According to the considered vector, one can get different values for the summation in eq. (10) and hence for the energy shift.
@@
Interestingly, this result can also be linked to the description of information processing properties of the out-of-equilibrium XXZ network with all-to-all connectivity analyzed in [5, 14]. There, an ideal-observer analysis of the power spectra had shown that the uncertainty with which the quantum network encodes the number of frequency peaks - and therefore the number of excitations - linearly scales with the number itself. In neuroscience, this is the content of Weber’s law describing the perception of the number of items in a space, or numerosity perception. This phenomenology is experimentally observed under very general conditions, no matter the perception channel (i.e. visual or auditory), nor the way the items are presented [72]. Importantly, neuroscience experiments indicate that the perception of number is interconnected to the perception of time and space intervals [77].
The overall bottom line is that even in the XXZ quantum network the emerging dynamical frequencies - or else the numerosity counting itself- comes from the number of consecutive gaps between the original levels. In fact, for small enough perturbations, anisotropy is found to just inject an uncertainty due to degeneracy lifting and the consequent spreading of the bands, the latter depending itself on the number of excitations introduced.
With this physical understanding at hand, we can now turn our attention to the transport properties of more general quantum networks, where we release either the constraint of constant coupling intensity of the links in the regular network, or the regular architecture itself by considering complex quantum networks. This is the subject of the next Section 3, where we explore the extent to which the space and time memory-preserving property of the regular XXZ network excitations can be engineered in more general networks; We will focus, in particular, to the case of one initial excitation or, more generally, in the m=L/2+1m=-L/2+1 total $z$-magnetization sector.
@@ 3 Transport properties in generalized quantum networks
So far we built a theoretical understanding about the out-of-equilibrium dynamical behavior of the closed quantum spin network governed by the XXZ Hamiltonian, for a network with all-to-all connectivity architecture of equally weighted links. For this case, we could resort to semi-analytical means. In this section, we start from the understanding developed in Section 2 to numerically investigate the role played by (i) the interaction graph in a complex quantum system architecture, and (ii) the range of the interactions according to a power law when all sites are connected between them. For simplicity, we stick to the case of one single excitation, having developed in Section 2 an intuition for multiple excitations: we will come back to this point in the concluding Section 4. As an example for complex quantum network, we choose the small-world architecture: besides being per se relevant for diverse problems, we are here interested to its property of well interpolating between the cases of regular and random graphs after tuning only two parameters.
We analyze the underlying physics by specializing our analysis to a localization quantifier called Inverse Participation Ratio (IPR), see Appendix B.
@@ 3.1 Small-world networks
A number of real-life networks relevant to diverse phenomena, ranging from biology to social sciences, are characterized by a high level of clustering and, at the same time, by a low average distance between the vertices: this is the so-called “small-world phenomenon” that illustrates e.g. social networks and the World Wide Web [78].
@@
Such a network can be built from a regular-lattice graph where each node is linked to its first $k$ neighbors, and then introducing a rewiring probability $\beta$ that one edge connects the starting node to a different one. The parameter $2k$ drives the level of network complexity, as it denotes the number of initial connections for each site. The parameter $\beta$ governs the level of graph’s randomness, bridging between the regular-lattice case with $\beta=0$ and the fully random graph with $\beta=1$. The all-to-all connected graph is obtained in the limit $\beta=0$ and 2k/(L1)=12k/(L-1)=1. The construction mechanism is illustrated in Fig. 4(a)(a).
@@
We complete this introductory section by recalling the main qualifiers and quantifiers for the networks architectures considered in complex-network theory, that will be used in the following. For details, we refer to the Appendix C. One first connectivity quantifier is the node degree, denoted by the word “deg”, that is the number of edges linking one node to others. The coupling range is quantified by the topological distance distij, counting the number of links connecting two sites in their shortest path: a global distance quantifier $\mathcal{D}$ can then be usefully introduced, as the average of the topological distance over all the pairs of nodes. Finally, the tendency to aggregate is quantified by the global clustering coefficient $\mathcal{C}$, that is the average proportion of total triangles (meaning closed loops of three vertices) within a given neighborhood, as related to the maximum number of allowed connections according to the local degree. All-to-all connected graphs are the limiting case of low topological distance, i.e. tending to 1, and high clustering coefficient, i.e. tending to 1; in the opposite limiting case of nearest-neighbor networks both local and global clustering are 0, since no triangles between nodes can be observed, while average distance linearly scales as the number of sites $L$.
3.2 Memory effects in small-world networks
Here we analyze the memory effect discussed in Section 2.1 in small-world networks, starting with a single excitation, i.e. a single spin oriented along $z$ while all the other spins are along the opposite direction. We use the same technique of Section 2.1 in which we take the Fourier transform of $S^{z}_{\text{initial}}(t)$. In presence of a single excitation only one high-energy resonance (whose energy scale with the system size) is present in the dynamics for all-to-all interaction, allowing us to count the number of excitations. Here we analyze how the Fourier spectrum changes as we decrease the connectivity. Interestingly, even if the number resonances increases as we decrease the connectivity, the high-energy resonances persists even for lower connectivity, see Fig. 4(c)(c). This result could allow to extend the notion of memory effect to generic small-word network. Indeed, it shows that signatures of the memory-preserving effect survive even for less connected network. Clearly, it would be interesting to study if the memory-preserving effect persists in small-world networks also in the many excitation case. This will be the topic of a future work.
Figure 4: Construction of small-world graphs and their memorization properties. $(a)$ Small-world network construction (see text). Top to bottom: a regular graph with $L=7$ sites and $k=2$ nearest-neighbor connections, and different rewired links (green and orange) with respect to the original graph, according to the value of the probability $\beta$. (b)(e)(b)-(e) Examples of intensity spectra for average $z$-magnetization at the initial site for one quenched excitation along the $z$-axis in four different networks with $L=33$ sites. $(b)$ All-to-all network, with delta-like peaks at frequencies $\omega^{*}=\pm JL/2$=33, see Eq. 4. $(c)$ Small-world, highly-connected graph with $(k,\beta)$=(15,0.5), possessing additional low-frequency components. $(d)$ Small-world, highly-connected graph with $(k,\beta)$=(13,0.5), having small high-frequency contributions. $(e)$ Sparsely connected small-world with $(k,\beta)$=(1,0.2), having no high-frequency peaks. The data has been detrended before the Fourier transform by removing the zero-frequency component.
@@ 3.3 Inverse Participation Ratio
The transport properties of one excitation in such networks can be characterized by means of a delocalization quantifier, the Inverse Participation Ratio (IPR)
@@
defined as the sum of the squared trace elements of the density matrix $\hat{\mathcal{\varrho}}$ over single-excitation localized states $\{|s\rangle\}=\{|\downarrow\downarrow\dots\downarrow\uparrow\downarrow\dots\downarrow\rangle\}$, where the $\uparrow$ in the previous Pauli string is a spin up (along $z$) located at site $s$; this basis has dimension $L$ and constitutes one of the subspaces of the total $z$-magnetization operator $\hat{\mathcal{M}}_{z}$ with which $\hat{\mathcal{H}}$ can be block-diagonalized. While we refer to Appendix B for a complete definition, we here highlight that by definition the IPR values range from $1/L$ for the complete delocalized state, i.e. infinite-temperature case, to $1$ for the perfect localization on a single eigenstate. If needed, generalizations of this quantifiers to the case of multiple excitations can be introduced, as in [4].
Before proceeding, it is useful to briefly discuss the form of the IPR function for a single excitation in the all-to-all case, as in Section 2: following the theoretical discussion in Section 2.1, one can easily derive the analytical function as the sum of a constant term and two cosine functions, oscillating with frequencies $\omega^{*}$ and $2\omega^{*}$, as $\omega^{*}$ is the same frequency of the local $z$-magnetization signals that is the energy gap between the levels $|l=L/2,m=-l/2+1\rangle$ and $|l=L/2-1,m=-L/2+1\rangle$, see Eq. 4. Therefore, in the following we will also show the IPR average for the all-to-all case over a temporal window $2\pi/\omega^{*}$. The way delocalization happens in small-world networks is strongly dependent on the initial conditions, which for our case translates to the site of the first excitation and its degree; this is the focus of Section 3.4.
3.4 Dependence on initial conditions
@@
We start by seeking an understanding of the importance of initial conditions, as we have done in Section 2. Since the analytical machinery that we established for regular networks seems to not apply for irregular graphs, we seek for alternatives. It turns out that this is possible at least for the short-time IPR behavior, by means of the so-called cascade model introduced in [79]: in essence, this monitors the initial state $|s(0)\rangle$ coupling with other basis elements at different moments $|s(t)\rangle$.
@@
The underlying idea is as follows, see Fig. 5(a)(a) for the concept. At $t=0$, only the subset of state vectors $\mathcal{F}_{0}$ containing $|s\rangle$ participates in the temporal evolution. Immediately after, to first-order approximation the temporal evolution operator connects the initial state to the basis vectors $|f_{i}\rangle$ for which the matrix elements $\langle f_{i}|\hat{\mathcal{H}}|s\rangle$ are non-vanishing, thus extending the subset of involved states to $\mathcal{F}_{1}$. More precisely, the subset of states grows as
@@
Figure 5: Understanding transport in small-world graphs at early times. $(a)$ Concept of the cascade model (see text). Top: an initial excitation (red circle) in a generic network (black edges) is located at site 1, as denoted by the state $|s=1\rangle=|\uparrow\downarrow\dots\downarrow\rangle$ and the red circle opposed to the blue ones. Bottom: after a short temporal evolution (black arrow) the initial excitation has spread (gradient-colored circles) across site 1’s first neighbors (orange edges), such that the state is a superposition of $|s=1\rangle$ and states $|f_{2}\rangle=|\downarrow\uparrow\downarrow\dots\downarrow\rangle$, $|f_{4}\rangle$, $|f_{5}\rangle$ and $|f_{7}\rangle$. (b)(c)(b)-(c) Comparison of the IPR’s time evolution as computed from the cascade model (orange curves) and from a numerical simulation (blue curves) for the following small-world networks: $(b)$ $k=1$, initial site with degree 2; $(c)$ $k=3$, initial site with degree 4.
@@
Borgonovi et al. have found a transparent expression for the evolution of IPR at short times, in the approximation where any back-flow to the initial site can be neglected [79]. This is:
@@
where the sum is performed over the first-neighbor sites of the initial one, and $\mathcal{N}_{1}$ is the number of such nodes plus one (referring to vertex $s$). Let us better understand Eq. 13: as the survival probability $P_{s}$ decreases because of the excitation spreading, the initial site couples with its first neighbors in the interaction topology. In other words, the IPR evolution is strictly dependent on the graph upon which it lives, through its adjacency matrix $A_{ij}$ describing the existing connections and being proportional to the XX Hamiltonian itself, when written in the restricted basis of computational vectors having one excitation. Notice that $A_{ij}=1$ if sites $i$ and $j$ are linked, else $A_{ij}=0$. Under a different perspective, the degree of the initial node (i.e. the number of first neighbors in the graph) governs the transport onset at short times.
@@
A summary of the analytical results is presented in Fig. 5. In Fig. 5(b)(c)(b)-(c), we compare the simulated (blue curves) IPR time behavior with the predictions of the cascade model (orange curves) for two complex network graphs that are sparsely $(b)$ and highly $(c)$ connected, respectively. For earlier times we find a good agreement between the IPR simulations and the predictions provided by the cascade model. In fact, this is able to predict the sharp deviation from the all-to-all case, heralding the breaking of the spectral degeneracy and the halt of cooperative shielding. The later times behavior from the cascade model misses the bumps that are likely due to the backflow processes, that are neglected in the model approximations.
Another illuminating way of looking at the impact of initial conditions on excitation spreading regards increasing connectivities close to the limit of all-to-all connected networks. The results are illustrated in Fig. 6(ab)(a-b), where we consider a network with a given 2k=L22k=L-2 parameter, closest to all-to-all case but not-fully connected, and impose a non-zero rewiring probability providing the sites with different degrees, i.e. different number of links with other elements. Then, we compute the IPR by picking different initial sites according to their degree. We do this for both long, Fig. 6 $(a)$, and short, Fig. 6 $(b)$ temporal windows, and we observe that the IPR time-evolution behavior for the all-to-all case soon breaks into the general small-world system behavior. Here, the degeneracy lifting due to the symmetry reduction with respect to the all-to-all case makes the IPR decreasing over time. The stationary value of the IPR critically depends on the degree of connectivity of the initial site.
We quantify the departure of small-world behavior from the all-to-all connectivity case in Fig. 6(c)(c), where we introduce an indicator for the difference the distance:
$\Delta_{\text{IPR}}(t)=|\text{IPR}_{\text{a.a.}}(t)-\text{IPR}_{\text{s.w.}}(t)|\,.$ (14)
[…]
Figure 6: Understanding transport in highly-connected networks. $(a)$ IPR behaviors for long periods of time for varying initial sites, denoted by their degree, in a fixed network with $L=33$, 2k=L22k=L-2, $\beta=0.1$. $(b)$ Same as $(a)$ but for short times; here $\omega^{*}=JL/2=33$ dictates the oscillation periods of the all-to-all connected IPR. $(c)$ IPR distance (14) between all-to-all and small-world networks graphs as a function of the initial site’s degree related to L1L-1 at increasing times as in the legend.
3.5 Infinite-time localization depending on graph properties